HIGH RESOLUTION INFRARED AND MICROWAVE SPECTRA OF NH_{3}-HCCH AND NH_{3}-OCS COM-PLEXES: STUDIES OF WEAK C-H⋯N HYDROGEN BOND AND ELECTRIC MULTIPOLE INTERAC-TIONS

\underline{XUNCHEN LIU}, YUNJIE XU, Department of Chemistry, University of Alberta, Edmonton, Canada, T6G 2G2.

C-H⋯N weak hydrogen bond is of much current interest. We report the first high resolution infrared spectroscopic study of a prototypical C-H⋯N bonded system, i.e. NH_{3}-HCCH, at the vicinity of the ν4 band of NH_{3}. The spectrum has been recorded using an infrared spectrometer equipped with an astigmatic multipass cell aligned for 366 passes and a room temperature external cavity quantum cascade laser at the 6 μm region. The perpendicular band spectrum of symmetric top rotor observed is consistent with the previous microwave^{a} and infrared^{b} studies at 3 μm. We also extended the previous microwave measurement to higher J and K. For the related NH_{3}-OCS complex, microwave spectrum of J up to 6 and infrared spectrum at the vicinity of the ν4 band of NH_{3} have been recorded and analyzed for the first time. Comparison has been made with the previously studied isoelectronic complexes such as NH_{3}-N_{2}O^{c} and NH_{3}-CO_{2}.^{d} The source of the difference will be discussed with the aid of ab initio calculations.

aG.T. Fraser, K.R. Leopold, and W. Klemperer, J. Chem. Phys. 80(4), 1423, (1984)

bG. Hilpert, G.T. Fraser, and A.S. Pine, J. Chem. Phys. 105(15), 6183, (1996)

cG.T. Fraser, D.D. Nelson, JR., G.J. Gerfen, and W. Klemperer, J. Chem. Phys. 83(11), 5442, (1985)

dG.T. Fraser, K.R. Leopold, and W. Klemperer, J. Chem. Phys. 81(6), 2577, (1984)